### v1

# Expert-Guided Diffusion Planner for Auto-bidding

Abstract.
Auto-bidding is extensively applied in advertising systems, serving a multitude of advertisers. Generative bidding is gradually gaining traction due to its robust planning capabilities and generalizability. In contrast to traditional reinforcement learning-based bidding, generative bidding does not rely on the Markov Decision Process (MDP) exhibiting superior planning capabilities in long-horizon scenarios. Conditional diffusion modeling approaches have demonstrated significant potential in the realm of auto-bidding. However, relying solely on return as the optimality condition is weak to guarantee the generation of genuinely optimal decision sequences, lacking personalized structural information. Moreover, diffusion models’ t-step autoregressive generation mechanism inherently carries timeliness risks. To address these issues, we propose a novel conditional diffusion modeling method based on expert trajectory guidance combined with a skip-step sampling strategy to enhance generation efficiency. We have validated the effectiveness of this approach through extensive offline experiments and achieved statistically significant results in online A/B testing, achieving an increase of 11.29% in conversion and a 12.35% in revenue compared with the baseline.
Auto Bidding, Generative Learning, Diffusion Model
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
The bidding mechanism of online advertising platforms, as a core component of modern digital marketing (Yuan et al., 2013; Wang et al., 2017; Adikari and Dutta, 2015; Evans, 2009), remains pivotal in enabling businesses to target audiences and drive commercial outcomes precisely. However, the inherent volatility of dynamic market conditions and the intricate diversity of user behaviors, combined with the need to process trillions of daily ad impressions in real time (Guo et al., 2024a; Johnson et al., 2020), expose dual challenges for traditional manual bidding strategies. Advertisers must constantly modify their bid prices to keep up with the constantly changing competitive bidding environment, which requires extensive knowledge in the field. Still, they frequently do not reach the best results.
To address these challenges, Ad platforms have developed auto-bidding services (Wen et al., 2022; Aggarwal et al., 2024; Lin et al., 2024) to simplify bid optimization for advertisers. Through carefully designed bidding strategies that account for various advertising factors (such as impression distributions, budgets, and cost constraints), these services automatically set bid prices for each impression opportunity. Various advanced technologies have been recently adopted to meet these demands, including Reinforcement Learning (RL)-based and Generative methods.
Reinforcement learning models the bidding process as a Markov Decision Process (MDP) and has been widely adopted in automated bidding systems, demonstrating consistent performance enhancements across diverse advertising scenarios (He et al., 2021; Lin et al., 2023; Mou et al., 2022). However, conventional reinforcement learning approaches grounded in Markov conditional independence assumptions and dynamic programming principles exhibit inherent limitations in bidding effectiveness, which primarily stems from (1) severe cumulative error propagation due to extended decision horizons in bidding sequences and (2) ignoring the impact of long-term dependencies in the bidding environment (Zhu et al., 2023).
Generative bidding demonstrates superior long-horizon planning capabilities and has gained significant traction.(Li et al., 2024) applies Decision Transformer to conduct sequential planning. (Guo et al., 2024b) applies a diffusion planner, incorporating influences from preceding states and enabling extended-horizon strategic planning.
However, applying conditional diffusion models to bidding faces two key challenges. First, existing approaches condition cumulative returns as a single scalar hyperparameter during inference, leading to weak guidance due to (i) lack of personalization and (ii) excessively large solution space satisfying the scalar constraint, thereby hindering convergence to optimal solutions. Second, the online bidding module’s strict latency requirements conflict with the T-step iterative sampling process inherent to diffusion models.
To tackle the challenges previously discussed, we introduce a Expert-Guided Diffusion planner for auto-Bidding (EGDB). Firstly, we propose a blended-forcing mechanism that synergistically integrates teacher-forcing stabilization and VAE-enhanced decode-forcing. It enables precise imitation of expert trajectories while proactively exploring latent optimal bidding patterns beyond observed demonstrations. Secondly, building upon the refined expert guidance, we devise a dual-conditioned planner that implicitly incorporates expert behavioral semantics while explicitly enforcing domain constraints (e.g., budget), achieving principled adaptation to dynamic auction environments. Finally, to overcome the inherent latency of iterative diffusion, we develop a step-skipping strategy that reduces inference steps without compromising action generation quality - a crucial advancement for real-time bidding systems. We demonstrate how this framework achieves state-of-the-art offline and online performance through comprehensive experiments.
In conclusion, our contributions are depicted as follows:
  • We explicitly construct expert demonstrations and leverage them to provide personalized guidance for the diffusion generation process.
  • To ensure training-inference consistency, we introduce a Variational Auto-Encoder (VAE) to model expert inherent behaviors, coupled with blended-forcing techniques to mitigate error accumulation and enhance training efficiency.
  • We adopt accelerated sampling with skip-step to further speed up the generation process.
  • To verify the effectiveness of the proposed method, we conduct extensive offline experiments and perform ablation studies to assess the contribution of each component. Moreover, we demonstrate its practical efficacy through online deployment.
2. Related Work
Offline Reinforcement Learning. In recent years, Offline reinforcement learning (Offline RL) has emerged as a pivotal paradigm for decision-making. It aims to learn optimal policies solely from static datasets of historical interactions, without requiring online environment exploration. This paradigm addresses scenarios where allowing partially trained agents to interact with the environment could lead to safety risks or substantial costs. Offline RL methods (Kostrikov et al., 2022; Fujimoto et al., 2019; Kumar et al., 2020; Chen et al., 2021) address two critical challenges—distribution shift and extrapolation errors-through distinct methodological innovations. Implicit Q-Learning (IQL) (Kostrikov et al., 2022) tackles distribution shift by implicitly estimating optimal actions through upper expectile regression on state values, suppressing extrapolation errors via asymmetric Q-updates that prioritize in-distribution high-value actions. Batch-Constrained Q-learning (BCQ) (Fujimoto et al., 2019) combines a variational autoencoder (VAE) with Q-filtering to restrict policy updates to data-supported regions, minimizing overestimation risks. Conservative Q-Learning (CQL) (Kumar et al., 2020) explicitly regularizes Q-values with theoretical guarantees, penalizing out-of-distribution (OOD) actions through conservative value bounds. In contrast, Decision Transformer (DT) (Chen et al., 2021) circumvents value functions entirely by reframing RL as sequence modeling, autoregressively generating actions conditioned on historical trajectories to eliminate bootstrapping errors and OOD queries.
Diffusion Models. Through their unique stochastic denoising mechanisms, diffusion models (Ho et al., 2020; Song et al., 2020) have established themselves as a dominant paradigm in image synthesis, capable of generating high-fidelity data via Markov chain-based iterative refinement. These models can be divided into two categories: conditional diffusion and unconditional diffusion. Conditional diffusion models (Fu et al., 2024) incorporate additional information (e.g., class labels, text prompts, or environmental states) to steer the generation process, enabling targeted synthesis. While unconditional diffusion models (Heng et al., 2024) generate data from pure noise without external guidance, excelling in capturing the intrinsic distribution of the training data. Owing to the powerful capabilities of diffusion models, they have been widely applied across various fields, particularly in decision-making. (Hansen-Estruch et al., 2023) proposes IDQL, a reinforcement learning framework that integrates implicit Q-learning with diffusion policies in an actor-critic architecture. (Wang et al., 2022) leverages diffusion models to capture complex, multimodal behavior distributions from static datasets. (Chen et al., 2022) uses diffusion models to more accurately fit the complex distribution of behavioral policies and avoids selecting unseen actions that are out of the data distribution during training.
Auto-bidding. Auto-bidding, which is primarily used in online advertising auction scenarios, dynamically adjusts advertisers’ bidding strategies to maximize advertising effectiveness (such as clicks, conversions, or sales) while meeting constraints like budgets and return on ad spend (ROAS). RL-based methods are widely used in auto-bidding because they dynamically optimize bids in complex advertising environments and balance exploration and exploitation. (Cai et al., 2017) proposes a reinforcement learning-based framework to optimize real-time bidding strategies in display advertising. (He et al., 2021) proposes a generalized framework for optimizing real-time bidding strategies under diverse budget and key performance indicator (KPI) constraints, dynamically adjusting bid parameters via reinforcement learning to maximize campaign value while ensuring compliance with advertiser requirements. (Jin et al., 2018) explores multi-agent reinforcement learning for real-time bidding in display advertising. Recent studies (Gao et al., 2025; Guo et al., 2024b) have increasingly explored generative auto-bidding strategies in online advertising. (Gao et al., 2025) leverages Decision Transformer (Chen et al., 2021) to model sequential dependencies and historical auction dynamics, enabling adaptive bidding strategy optimization by autoregressively generating context-aware bid adjustments aligned with long-term advertiser objectives. (Guo et al., 2024b) directly models the correlation between the return and the entire trajectory and uses a conditional diffusion for bid generation.
3. Preliminary
In this section, we outlined the general bidding and auction tasks framework and examined the architecture of generative diffusion models. To prevent any confusion regarding subscripts in the subsequent discussion, it should be noted that $k$ denotes the diffusion timesteps, while $t$ represents the position of the trajectory $\tau$.
3.1. Formulation of Bidding Problem
In online advertising auctions, the bidding problem typically involves maximizing a specified objective while adhering to certain constraints. For each advertiser $i$, the most frequently examined are budget ($B_{i}$) and return-on-spend (RoS) constraints ($C_{i}$). Consequently, the bidding problem can be formulated as follows (He et al., 2021):
(1) $\begin{aligned} \max\sum_{j}\mathbb{A}_{ij}\cdot v_{ij}\quad s.t.\sum_{j}(\mathbb{A}_{ij}\cdot b_{j})&\leq B_{i},\\ \frac{\sum_{j}(\mathbb{A}_{ij}\cdot\mathbb{P}_{ij})}{\sum_{j}(\mathbb{A}_{ij}\cdot v_{ij})}&\leq C_{i}.\end{aligned}$
Here, $b_{j}$ denotes the bids submitted by all advertisers for each auction $j$, and $v_{ij}\in R^{+}$ signifies the revenue generated by the $i$-th advertiser through auction $j$. The term $\mathbb{A}_{ij}(b_{ij})\in[0,1]$ indicates whether advertiser $i$ wins auction $j$, and $\mathbb{P}_{ij}$ represents the payment made by advertiser $i$ to secure auction $j$.
Figure 1. The foundational framework of our proposal consists of the following components: (1) the block of expert-guided conditional diffusion (EGCD) and (2) the overall architecture of  EGDB.
3.2. Diffusion Model
The diffusion model (DM) (Ho et al., 2020) functions via forward and reverse processes, and it is widely used in recommender systems. Starting with an initial data sample $x_{0}\sim q(x_{0})$, the forward process incrementally introduces Gaussian noise through a Markov chain over $K$ steps, resulting in latent variables $x_{1:K}$. The forward transition is defined by $q(x_{k}|\boldsymbol{x}_{k-1})=\mathcal{N}(\boldsymbol{x}_{k};\sqrt{1-\beta_{k}}\boldsymbol{x}_{k-1},\beta_{k}\boldsymbol{I})$, where $k\in\{1,\cdots,K\}$ denotes the diffusion step, $\mathcal{N}$ represents the Gaussian distribution, and $\beta_{k}\in(0,1)$ controls the noise intensity at step $k$. As $K$ approaches infinity, $x_{K}$ converges to a standard Gaussian distribution. The DM reconstructs $x_{k-1}$ from $x_{k}$ by learning a denoising trajectory in the reverse process. Commencing from $x_{K}$, it progressively approximates $x_{k}\to\boldsymbol{x}_{k-1}$ through $p_{\theta}(\boldsymbol{x}_{k-1}|\boldsymbol{x}_{k})=\mathcal{N}(\boldsymbol{x}_{k-1};\mu_{\theta}(\boldsymbol{x}_{k},k),\Sigma_{\theta}(\boldsymbol{x}_{k},k))$, where $\mu_{\theta}(x_{k},k)$ and $\Sigma_{\theta}(x_{k},k)$ are outputs of a neural network parameterized by $\theta$, capturing the intricate generative dynamics.
3.3. Classifier-free Diffusion Guidance
The sequential nature of auto-bidding requires capturing both temporal dependencies and constraint satisfaction (e.g., budget limits). Traditional autoregressive methods often suffer from error accumulation in multi-step generation, while standard generative models struggle to incorporate dynamic constraints. Consequently, following (Guo et al., 2024a), we formulate the sequential decision problem through conditional generative modeling:
(2) $\max_{\theta}\mathbb{E}_{\tau\sim D}\left[\log p_{\theta}\left(x_{0}(\tau)\mid y(\tau)\right)\right]$
[…]
(3) $q\left(x_{k+1}(\tau)\mid x_{k}(\tau)\right),\quad p_{\theta}\left(x_{k-1}(\tau)\mid x_{k}(\tau),y(\tau)\right),$
where $q$ represents the forward process in which noises are gradually added to the trajectory, while $p_{\theta}$ is the reverse process where a model is used for denoising.
We model the forward process $q\left(x_{k+1}(\tau)\mid x_{k}(\tau)\right)$ via diffusion over states, where:
(4) $x_{k}(\tau):=\left(s_{1},\dots,s_{t},\dots,s_{T}\right)_{k},$
where $s_{t}$ is a one-dimensional vector, including real-time bidding context, and $x_{k}(\tau)$ is a noise sequence of states which can be represented by a two-dimensional array where the first dimension is the time periods and the second dimension is the state values. Given $x_{k}(\tau)$, the forward process of the diffusion model can be represented as follows:
(5) $q\left(x_{k}(\tau)\mid x_{k-1}(\tau)\right)=\mathcal{N}\left(x_{k}(\tau);\sqrt{1-\beta_{k}}x_{k-1}(\tau),\beta_{k}\mathbf{I}\right),$
when $k\to\infty$, $x_{k}(\tau)$ approaches a sequence of standard Gaussian distribution, where we can make sampling through the reparameterization trick and then gradually denoise the trajectory to produce the final state sequence.
The classifier-guided diffusion model requires an accurate estimation of the guidance gradient based on the trajectory classifier $y(\tau)$, which may not be feasible. Therefore, we use a classifier-free guidance strategy to guide the generation. The core of Classifier-Free Guidance lies in replacing explicit classifiers with an implicit classifier, eliminating the need to compute explicit classifiers and their gradients directly. Following (Ajay et al., 2022; Ho and Salimans, 2022), who pioneered conditional guidance in diffusion models, we utilize a classifier-free guidance strategy with low-temperature sampling. During the training phase, we need to train the unconditional model $\epsilon_{\theta}\left(x_{k}(\tau),k\right)$ and the conditional model $\epsilon_{\theta}\left(x_{k}(\tau),y(\tau),k\right)$. However, these can be represented by a single shared model by randomly dropping the condition. The final output is obtained through a linear extrapolation of the conditional and unconditional generation predictions:
(6) $\hat{\epsilon}_{k}:=\epsilon_{\theta}\left(x_{k}(\tau),k\right)+\omega\left(\epsilon_{\theta}\left(x_{k}(\tau),y(\tau),k\right)-\epsilon_{\theta}\left(x_{k}(\tau),k\right)\right),$
where the guidance coefficient $\omega$ can be adjusted to control the trade-off between the realism and diversity of generated samples. After that, we can sample from $p_{\theta}\left(x_{k-1}(\tau)\mid x_{k}(\tau),y(\tau)\right)$ as follows:
(7) $x_{k-1}(\tau)\sim\mathcal{N}\left(x_{k-1}(\tau)\mid\mu_{\theta}\left(x_{k}(\tau),y(\tau),k\right),\Sigma_{\theta}\left(x_{k}(\tau),k\right)\right),$
When serving at period $t$, we first sample from an initial trajectory $x^{\prime}_{k}(\tau)\sim\mathcal{N}(0,\mathbf{I})$ and assigns the history states $s_{0:t}$ into it. Then, we can sample predicted states with the reverse process recursively by
(8) $x^{\prime}_{k-1}(\tau)=\mu_{\theta}\left(x^{\prime}_{k}(\tau),y(\tau),k\right)+\sqrt{\beta_{k}}z,$
where $z\sim\mathcal{N}(0,\mathbf{I})$. Given $x^{\prime}_{0}(\tau)$, we can extract the next predicted state $s^{\prime}_{t+1}$, and determine how much it should bid to achieve that state. In this setting, we apply the widely used inverse dynamics (Agrawal et al., 2016) with a non-Markov state sequence to determine current bidding parameters at period $t$:
(9) $\hat{a}_{t}=f_{\phi}\left(s_{t-h:t},s^{\prime}_{t+1}\right),$
where $\hat{a}_{t}$ contains predicted bidding parameters at time $t$ and $h$ is the length of history states.
4. Methodology
In this section, we present a comprehensive description of the proposed Expert-Guided Diffusion planner for auto-Bidding ( EGDB ) framework. We begin by formalizing the core mechanism of Blended Forcing with Optimal Expert in Section 4.1. Next, in Section 4.2, we elaborate on the Expert-guided Conditional Planner that translates blended signals into structured action sequences. Finally, Section 4.3 details the Efficient Action Generation module designed for real-time execution. The overall architecture is illustrated in Figure 1.
4.1. Blended Forcing with Optimal Expert
To address the bidding problem outlined in Equation 1, we initially aim to generate expert trajectories. To concentrate on a single bidding agent, we remove the subscript $i$, assuming all other agents are fixed. We transform the linear programming (LP) function into its dual form. Assuming the auction is truthful, the bidding formula will produce an outcome corresponding to the optimal primal solution $b_{j}$ (Aggarwal et al., 2024). For each query $j$:
(10) $\mathbf{x}^{e}(\tau)=b_{j}:=\frac{1+\alpha_{c}\cdot C_{i}}{\alpha_{b}+\alpha_{c}}\cdot v_{j},$
where $\alpha_{b}$ represents the influence of budget constraints and $\alpha_{c}$ denotes the impact of constraints of cost.
Then, we can refine the $\mathbf{x}^{e}(\tau)$ to fully capture the latent expert behaviors through the blended-forcing mechanism in the training phase. As shown in Figure 1, it comprises two operations: (a) the teacher-forcing technique ($TF(\cdot)$) to imitate the expert behaviors more stably, and (b) the decode-forcing mechanism ($DF(\cdot)$) to explore more optimal expert behaviors through an extra Variational Auto-Encoder(Kingma, 2013) (VAE).
Teacher Forcing (TF), initially introduced in Recurrent Neural Networks (RNNs) (Williams and Zipser, 1989), is a training technique where the ground truth sequence is fed as input to the model at each timestep, rather than using its previous predictions. This approach effectively mitigates error accumulation in autoregressive models by aligning training dynamics with the true data distribution. However, a critical limitation arises during inference: the absence of ground truth forces the model to rely entirely on its own (potentially erroneous) predictions, leading to a mismatch between training and inference input distributions. This phenomenon, termed Exposure Bias, often results in degraded performance in real-world scenarios. Inspired by this technique, we introduce TF and use expert trajectories serving as ground truth sequences to help the model imitate expert behaviors.
While TF alleviates error propagation during training, its reliance on ground truth introduces a fundamental inconsistency between the training and inference phases. In complex dynamic environments (e.g., real-time bidding), traditional mitigation strategies like Scheduled Sampling (Bengio et al., 2015) — which heuristically blends ground truth and predicted tokens — fail to capture rapidly shifting environmental dynamics adaptively. The rigidity of such handcrafted mixing strategies often leads to suboptimal policy generalization, particularly when expert-guided constraints (e.g., budget allocation rules) must be preserved.
To bridge this gap, we propose a Variational Autoencoder (VAE)-enhanced hybrid mechanism called Blended Forcing, which implicitly models environmental dynamics through the lens of expert historical behaviors. The VAE’s latent space is trained to encode the distribution of optimal expert actions conditioned on environmental states. During inference, the VAE reconstructs a pseudo ground-truth action that approximates expert-level decisions given the current state, thereby serving as a dynamically adaptive substitute for the missing ground truth.
Moreover, we apply for coverage $\mathbf{x}^{e}(\tau)$ to $\hat{\mathbf{x}}^{e}(\tau)$ with a hyperparameter $\delta$ to harmonize the representation of expert, which can be formulated as:
(11) $\hat{\mathbf{x}}^{e}(\tau)=\mathbb{I}[TF(\mathbf{x}^{e}(\tau))_{\delta},DF(\mathbf{x}^{e}(\tau))_{1-\delta}],$
where $\mathbb{I}[\cdot]$ represents the indicator function, and in each $k$-th diffusion step, it will choose $TF(\cdot)$ in the probability of $\delta$ and $DF(\cdot)$ otherwise. As for the structure of $DF(\cdot)$, we implement one layer DNN for both the encoder and decoder with the learning parameters $\phi$ for reconstruction.
4.2. Expert-guided Conditional Planner
After we obtain the $\hat{\mathbf{x}}^{e}(\tau)$ as the extra condition for diffusion training, it serves as the implicit signal during the inference phase. Consequently, the overall framework should be guided more explicitly. Therefore, we incorporate the expected constraints $\mathbb{C}^{g}$ defined in Section 3.1 to formulate the overall condition $\mathbb{C}$ in a dual fashion, which can be depicted as:
(12) $\mathbb{C}=\underbrace{\hat{\mathbf{x}}^{e}(\tau)}_{IMPLICIT}\oplus\quad\underbrace{[f_{\text{}}(R(\tau)\oplus f_{\text{}}^{\prime}(C(\tau))]}_{EXPLICIT(\mathbb{C}^{g})}.$
Note that $R(\tau)$ denotes the total value the advertiser received as the return $R(\tau)=\sum_{t=1}^{T}r_{t}$. $f$ is min-max normalization, $f^{\prime}$ denotes a indicator function transfers constraint to $[0,1]$ corresponding to (Guo et al., 2024b) , and $\oplus$ represents the concatenation operation. Since DMs are optimized by maximizing the Evidence Lower Bound (ELBO) of the likelihood of observed input data (Ho et al., 2020), we adapt the optimization objective $\mathcal{L}_{\text{ddpm}}$ stated in Section 3.2 previously as:
(13) $D_{\text{KL}}(q(\mathbf{x}_{k-1}(\tau)\mid\mathbf{x}_{k}(\tau),\mathbf{x}_{0}(\tau))\mid\mid p_{\theta}(\mathbf{x}_{k-1}(\tau)\mid\mathbf{x}_{k}(\tau),\mathbb{C})).$
In line with prior researches  (Li et al., 2023; Yang et al., 2024), we utilize an alternative parameterization trick that predicts the target $\mathbf{x}_{0}$ rather than the added Gaussian noise $\epsilon$, to circumvent the representation collapse issue  (Wang et al., 2024) in discrete space as:
(14) $\tilde{\boldsymbol{\mu}}_{\theta}(\mathbf{x}_{k}(\tau),\mathbb{C},k)=\sqrt{\bar{\alpha}_{k-1}}f_{\theta}(\mathbf{x}_{k}(\tau),\mathbb{C},k)+\frac{(1-\bar{\alpha}_{k-1})\sqrt{\alpha_{k}}}{\sqrt{1-\bar{\alpha}_{k}}}\boldsymbol{\epsilon}.$
Drawing inspiration from the success of sequential recommendation models in  (Huang et al., 2024), we designed a novel block named Expert-guided Conditional Diffusion (EGCD) for the diffusion model architecture. This model is specifically built on the Transformer architecture (Vaswani et al., 2017), where we have substituted the traditional self-attention mechanism with cross-attention to leverage the conditional information fully.
Algorithm 1 Training with EGDB
1:Inputs: diffusion step $K$, expert trajectory $\boldsymbol{x}^{e}(\tau)$.
2:Parameters: $\Theta=\{\phi,\theta,\psi\}$
3:repeat
4:  Attain $\hat{\mathbf{x}}^{e}(\tau)$ through Eq. (11) with probability $\delta$
5:  Calculate $\mathbb{C}$ through Eq. (12)
6:  Sample $k\sim\text{Uniform}\{1,\cdots,K\}$ and $\boldsymbol{x}_{k}(\tau)[:t]\leftarrow\mathbf{s}_{0:t}$
7:  Estimate $\hat{\textbf{x}}_{0}(\tau)\leftarrow(\tilde{\mu}_{k-1},\Sigma_{k-1})$ through Eq. (13)
8:  $\boldsymbol{x}_{k-1}(\tau)\sim\mathcal{N}(\tilde{\mu}_{k-1},\alpha\Sigma_{k-1})$
9:  Take gradient descent step $\nabla_{\Theta}(\mathcal{L}_{\text{total}})$ through Eq. (19)
10:until converged
To formulate the query for cross-attention, we concatenate the original sequence generated in a single step, denoted as $x_{k}(\tau)$, the total number of denoising steps $k$, and the relevant conditions $\mathbb{C}^{g}$ as query.
(15) $\mathbb{O}=x_{k}(\tau)\oplus k\oplus\mathbb{C}^{g},\quad\mathbb{Q}=\text{LN}\left[\text{MLP}[\mathbb{O}]\right]$
It is noted that $\text{LN}(\cdot)$ and $\text{MLP}(\cdot)$ represent the layer normalization and dense layer, respectively, and cross-attention is defined as the process of extracting effective information from the expert sequence as follows.
(16) $\mathrm{Cross\_Attn}=\mathrm{Softmax}\left(\frac{(\mathbb{Q}W^{Q})(\mathbb{C}W^{K})^{T}}{\sqrt{d}}\right)(\mathbb{C}W^{V}),$
the Key and Value of Cross-Attention is defined in (12) and the parameter matrices $W^{Q}$$W^{K}$$W^{V}$ represent the weights associated with the transformations of the original $Q$,$K$, and $V$, and $d$ denotes the dimension after the vectors have been concatenated. The latent representation $\mathbb{H}_{0}$ can be obtained from cross-attention, and $x_{k-1}(\tau)$ is obtained as follows.
(17) $\mathbb{H}_{1}=\text{FFN}\left[\text{LN}[\mathbb{H}_{0}+\mathbb{O}]\right],\ \text{and}\ x_{k-1}(\tau)=\text{MLP}\left(\mathbb{H}_{1}+\mathbb{H}_{0}\right),$
where FFN is Feed Forward Network, and $\mathbb{H}_{0}$ and $\mathbb{H}_{1}$ are add as output.
4.3. Efficient Action Generation
4.3.1. Accelerate Sampling.
By employing Equation 14, the inference Markov transition of  EGDB  can be defined as $p_{\theta}(\mathbf{x}_{k-1}(\tau)\mid\mathbf{x}_{k}(\tau))$. Each step is designed to predict $\hat{\mathbf{x}}_{0}(\tau)$ and provide $\mathbf{x}_{k-1}(\tau)$, which is less noisy than $\mathbf{x}_{k}(\tau)$ for the next step. The time cost of the inference process increases linearly with the total number of steps. A natural approach to accelerate this process is to skip several intermediate steps if the early approximations are accurate enough and require no further adjustment. Here, we define $p_{\theta}(\mathbf{x}_{k-\gamma}(\tau)\mid\mathbf{x}_{k}(\tau))\rightarrow p_{\theta}(k-\gamma\mid k)$ with the number of skipped steps $\gamma$ as:
(18) $p_{\theta}(k-\gamma\mid k)=\mathcal{N}(\mathbf{x}_{k-\gamma}(\tau);\tilde{\mu}_{\theta}(\hat{\mathbf{x}}_{0}(\tau),t),\Sigma_{\theta}(\hat{\mathbf{x}}_{0}(\tau),t)),$
where the $\gamma$ governs the number of skip steps. This acceleration method significantly reduces the time overhead without sacrificing excessive accuracy. We will discuss the details in Section 5.2. After we get $\hat{\mathbf{x}}_{0}(\tau)$, we can generate the corresponding action $\hat{\mathbf{a}}_{t}$ through an inverse network $f_{\psi}(\cdot)$.
Algorithm 2 Generating Optimal Actions
1:Inputs: diffusion step $K$, expert trajectory $\mathbf{x}^{e}(\tau)$.
2:Parameters: $\Theta=\{\phi,\theta,\psi\}$
3: Get history of states $\mathbf{s}_{0:t}$;
4: Sample $\mathbf{x}_{K}(\tau)\sim\mathcal{N}(0,\beta_{K}I)$, and $\hat{\mathbf{x}}^{e}(\tau)\sim\text{Norm}[0,1]$
5: Decode $\hat{\mathbf{x}}^{e}(\tau)$ with $z$ and calculate $\mathbb{C}$ through Eq. (12)
6:repeat
7:for $k=\{K,K-1,\cdots,1\}$ do
8:  $\mathbf{x}_{k}(\tau)[:t]\leftarrow\mathbf{s}_{0:t}$
9:  Estimate $\hat{\textbf{x}}_{0}(\tau)\leftarrow(\mu_{k-\gamma},\Sigma_{k-\gamma})$ through Eq. (18)
10:  $\mathbf{x}_{k-1}(\tau)\sim\mathcal{N}(\tilde{\mu}_{k-1},\alpha\Sigma_{k-1})$
11:end for
12: Extract $(\mathbf{s}_{t-h:t},\mathbf{s}^{\prime}_{t+1})$ from $\hat{\mathbf{x}}_{0}(\tau)$
13: Generate $\hat{\mathbf{a}}_{t}=f_{\psi}(\mathbf{s}_{t-h:t},\mathbf{s}^{\prime}_{t+1})$;
4.3.2. Joint Optimization.
As for optimization, considering that VAE is employed for extracting information from expert sequences, the VAE loss (Kingma, 2013) must also be incorporated as $\mathcal{L}_{\text{exp}}$, which primarily consists of the reconstruction loss of $z$ and the constraint that the intermediate latent variable $z$ follows a normal distribution. Moreover, to inverse actions through the series of states generated by the conditional diffusion planner, we utilize the inverse dynamics loss $\mathcal{L}_{\text{inv}}$ to optimize the reconstruction of the action patterns following (Guo et al., 2024b). Therefore, the optimized objective can be simplified as:
[…]
where the $\xi$ governs the reconstruction of the expert trajectory and action generation. The overall training and inference phases are shown in Algorithm 1 and 2, respectively.
5. Experiments
5.1. Experimental Setting
We conduct extensive experiments in offline environments to demonstrate the effectiveness of our proposed method. Four research questions are investigated in the following experiments:
  • RQ1: How does EGDB  perform compared to state-of-the-art auto-bidding baselines?
  • RQ2: How do the proposed components in EGDB  contribute to the final bidding performance?
  • RQ3: How do parameter settings impact the model’s performance?
  • RQ4: How does  EGDB  perform in real-world business scenarios compared to baselines, and is  EGDB  effective in improving inference efficiency?
5.1.1. Dataset & Environment.
We utilize Alibaba [^1] https://github.com/alimama-tech/AuctionNet simulated environment from the NeurIPS 2024 competition (Tianchi, 2024). The dataset covers 21 periods (7-27), each with over 500,000 impression opportunities and 48 steps. Each opportunity involves 48 bidding agents. The dataset contains over 500 million records, each with predicted conversion values, bids, auction details, and impression results. For a fair comparison, data from periods 7-13 is for training and 14-20 for testing.
5.1.2. Metrics.
To simplify the experimental evaluation, we kept the bids of 47 of the 48 advertisers unchanged during the test, only adjusted the bid of the remaining advertiser, and simulated the corresponding rewards and impressions.To evaluate these methods, we use the cumulative conversions of all advertisers, adjusted for penalties incurred by exceeding the Cost Per Action (CPA) ratio, as our metric.
(20) $\begin{aligned} score=\mathbb{P}(CPA,C)\sum_{i}x_{i}E_{i}V_{i},\quad\mathbb{P}(CPA,C)=\text{min}\{(\frac{CPA}{C})^{\lambda},1\}.\end{aligned}$
$x_{i}$ indicates whether the agent secures the impression, $E_{i}$ signifies if the ad is successfully shown to the user after winning, and $V_{i}$ denotes whether a conversion occurs after the user is exposed to the advertisement. The advertiser sets the target $CPA$, while $C$ represents the actual $CPA$ after completing the process. The penalty function defines a hyperparameter $\lambda$, usually set to 2.
5.1.3. Baselines.
To verify the effectiveness of our proposed method, we list some common reinforcement learning methods as benchmarks for comparison.
  • CQL (Kumar et al., 2020) addresses overestimation of out-of-distribution actions by introducing conservative constraints in the Q-function
  • IQL (Kostrikov et al., 2022) focuses on in-sample actions and avoiding the querying of values for unseen actions
  • BC (Torabi et al., 2018) is trained through supervised learning by directly replicating expert trajectory data
  • DT (Chen et al., 2021) Transformer architecture-based algorithm that maps trajectory data to action outputs through sequence modeling
  • DiffBid (Guo et al., 2024a) conditional diffusion modeling approach for bid generation
5.2. Experimental Results
5.2.1. Overall Performances (RQ1)
We compare the overall performance of  EGDB  with various baselines, as reported in Table 1. Our method consistently outperforms other approaches across all dates and achieves the highest scores on each date, validating the effectiveness of our proposed approach. This superiority arises from the robust generative capabilities of the diffusion model and the guidance provided by expert samples, which enhances the model’s ability to generate superior bidding trajectories. The IQL and CQL methods outperform BC, suggesting that traditional reinforcement learning can develop more effective strategies than simple supervised learning. Additionally, DT demonstrates superior performance compared to traditional reinforcement learning methods at certain times, indicating that taking sequence dependency into account is beneficial for optimizing bidding strategies.
Table 1. Performance comparison between baselines and  EGDB. The best performance of each column is bold, while the second-best is underlined.
Method P14 P15 P16 P17 P18 P19 P20
CQL 31.09 29.99 29.88 29.73 27.97 33.45 27.99
IQL 32.75 25.43 27.05 35.09 30.38 34.07 28.39
BC 28.53 27.37 30.55 28.64 29.13 31.70 26.43
DT 30.01 29.79 30.31 31.92 29.52 34.42 29.70
DiffBid 31.95 28.62 30.65 35.59 27.72 34.36 29.79
EGDB 38.91 33.69 37.56 39.51 34.73 38.66 34.01
5.2.2. Ablation Studies (RQ2)
To further clarify the contributions of each module in  EGDB , we conduct ablation experiments by evaluating various configurations of  EGDB . Table  2 presents the overall results of our ablation studies on the proposed components, revealing three key insights: (1) All modules contribute positively, with the CA module demonstrating a significant impact; (2) An adaptive fast sampling strategy can achieve optimal performance-accuracy trade-offs; (3) Blended forcing partially emulates expert sequence behaviors, indicating promising regularization potential.
Table 2. The ablation study results for our proposed  EGDB: (1) w/o Acc. $\rightarrow$ the model denoises step by step to replace fast-sampling; (2) w/o BF. $\rightarrow$ blended-forcing is removed and only the auto-encoder; (3) w/o CA. $\rightarrow$ Cross-Attention block is removed, and self-attention is adopted.
Method P14 P15 P16 P17 P18 P19 P20
w/o Acc. 37.52 31.83 36.40 37.89 33.66 36.89 32.92
w/o BF. 38.09 32.31 36.58 38.40 33.13 36.76 32.81
w/o CA. 35.29 29.54 31.29 35.49 31.91 35.72 31.08
All 38.91 33.69 37.56 39.51 34.73 38.66 34.01
5.2.3. Parameter Sensitivity (RQ3)
We further analyze the sensitivity of hyperparameter settings in  EGDB. As shown in Figure 3, we select the periods with the lowest (P15) and highest (P17) metrics for display to ensure our samples are representative.
  • $\gamma$ : The number of skip steps in accelerated sampling.
  • $\xi$ : balances the weight of the reconstruction of the expert trajectory and action generation loss within the loss function;
  • $\delta$ : A hyperparameter to harmonize the representation of the expert.
The experimental results indicate that the parameters $\gamma$ and $\xi$ are significantly influenced by hyperparameter changes, whereas $\delta$ is relatively less affected. Furthermore, it has been demonstrated that employing an appropriate skipping strategy can yield optimal results. Since a larger value of $\xi$ determines the model’s ability to reconstruct the expert and the action, placing greater emphasis on reconstruction may restrict the generalization capabilities of the Denoising Diffusion Probabilistic Model (DDPM), resulting in a gradual decline in performance metrics.
Figure 2. We examine the sensitivity of three key hyperparameters in the period of P15 (left) and P17 (right). The horizontal axis displays $\gamma$ for $\{1,2,4,8,16\}$, $\delta$ for $\{0,0.2,0.4,0.6,0.8\}$, and $\xi$ for $\{0.25,0.50,1.00,2.00,4.00\}$.
5.3. Online A/B Testing(RQ4)
Figure 3. The workflow of proposed  EGDB  in online system.
To verify the performance of the proposed method in real-world business scenarios, we selected a batch of the lowest-cost campaigns aimed at maximizing conversions within budget constraints. Diffbid (Guo et al., 2024a) serves as a baseline to evaluate the effectiveness of our proposed method, as our modifications are built upon it. Additionally, the performance of Diffbid on offline datasets is surpassed only by our proposed method. We conducted a budget-splitting and traffic-splitting online experiment from February 7, 2024, to February 13, 2024. The experiment results are presented in Table 3.
  • Feature: a series of window features with different window sizes, including the bid coefficient, remaining traffic, remaining budget, cumulative consumption, cumulative revenue, etc.
  • Action: the difference between two consecutive bid coefficients is represented as $a_{t}=b_{t+1}-b_{t}$, where $b_{t}$ denotes the bid coefficient for a specific advertiser at time $t$
Our method demonstrated a significant improvement compared to the baseline model. Specifically,  EGDB  increased conversions by 11.29% and revenue by 12.35%. In contrast to DiffBid, EGDB  primarily integrates Variational Autoencoder (VAE) expert trajectory generation and cross-attention-based denoising steps, which influence inference complexity. However, we utilized the skip sampling method to reduce 25% of the inference resources when serving the same advertising campaign while significantly maintaining the model’s effectiveness.
Table 3. Performance of  EGDB  in Online Testing.
Method Budget Campaign Conversions Revenue
Baseline (IQL) 269,686 247 265,155 1,188,336
Ours 269,686 247 295,090 1,335,214
compare +11.29% +12.36%
6. Conclusion
In this paper, we propose EGDB, an expert-guided conditional diffusion model for auto-bidding. To enhance the diffusion planner with a more robust and personalized condition, we extract expert trajectories and introduce the EGCD block, which integrates expert conditions into diffusion transformers. The inherent behavior of the expert is modeled using a Variational Autoencoder (VAE), and we propose a blended forcing technique to improve training efficiency. Additionally, we adopt accelerated sampling with a skip-step approach to speed up the generation process for matching online inference. Offline and online experiments have been designed to demonstrate superiority over baseline models. Future work will explore optimal path generation principles and hybrid architectures that combine diffusion models with the advantages of linear programming.